black pencils sell at two for 15 cents, and red pencils sell at 10 cents each. How many of each kind were sold if a total of 25 were sold for $2.00?
Answer by mananth(16949)

black pencils sell at two for 15 cents, and red pencils sell at 10 cents each. How many of each kind were sold if a total of 25 were sold for $2.00?
..
black pencils two for 15 cents,
so 1 for 7.5 cents
red pencils 10 cents each.
total of 25 were sold
for $2.00 = 200 cents
...
let black pencils be x numbers
red pencils will be 25-x
..
cost of black pencils = 7.5x
cost of red penci;s = 10(25-x)
..
7.5x+10(25-x)=200
7.5x+250-10x= 200
-2.5x=-250+200
-2.5x=-50
/-2.5
x= -50/-2.5
x=20 pencils ---- black
5 will be red
...
m.ananth@hotmail.ca
..